OpenStudy (anonymous):

the product (radical -8) (radical -2) is equal to a) 4i b) 4 c) -4 d) 3i radical2

OpenStudy (anonymous):

it's 4

OpenStudy (anonymous):

\[i \sqrt{8}*i \sqrt{2} = 2i \sqrt{2} * i \sqrt{2}= -2\sqrt{4}=-4\]

OpenStudy (anonymous):

doesn't work that way, look it up

OpenStudy (anonymous):

dockworker is right. thanks again :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

come again?

OpenStudy (anonymous):

c) -4 2 * 2 i^2 = -4